# Generative AI for Semantic Communication: Architecture, Challenges, and Outlook

### v2

@@ Abstract
Semantic communication (SemCom) is expected to be a core paradigm in future communication networks, yielding significant benefits in terms of spectrum resource saving and information interaction efficiency. However, the existing SemCom structure is limited by the lack of context-reasoning ability and background knowledge provisioning, which, therefore, motivates us to seek the potential of incorporating generative artificial intelligence (GAI) technologies with SemCom. Recognizing GAI’s powerful capability in automating and creating valuable, diverse, and personalized multimodal content, this article first highlights the principal characteristics of the combination of GAI and SemCom along with their pertinent benefits and challenges. To tackle these challenges, we further propose a novel GAI-assisted SemCom network (GAI-SCN) framework in a cloud-edge-mobile design. Specifically, by employing global and local GAI models, our GAI-SCN enables multimodal semantic content provisioning, semantic-level joint-source-channel coding, and AIGC acquisition to maximize the efficiency and reliability of semantic reasoning and resource utilization. Afterward, we present a detailed implementation workflow of GAI-SCN, followed by corresponding initial simulations for performance evaluation in comparison with two benchmarks. Finally, we discuss several open issues and offer feasible solutions to unlock the full potential of GAI-SCN.
@@ I Introduction
Recently, semantic communication (SemCom) is popularized as an emerging paradigm that promises to significantly alleviate the scarcity of communication resources in future wireless networks [1]. This is mainly benefited from prosper advancement in deep learning (DL) technologies that can drive semantic encoding and decoding models to achieve efficient and high-quality semantic refinement on desired meaning with low spectrum consumption. Moreover, through equipping both ends of the transceiver with equivalent background knowledge, i.e., provisioning massive data samples to serve diverse artificial intelligence (AI) learning and prediction tasks [2], the implicit meaning in conveyed content can be recovered with ultra-low semantic errors even under harsh channel conditions. However, realizing such superiorities obviously poses a huge demand on data acquisition for constructing background knowledge and pre-training DL-driven semantic models. Meanwhile, considering that the achievable semantic performance is essentially confined by the quality of pre-training data used, existing SemCom systems still lack sufficient context reasoning capabilities, i.e., accurate semantic calibration and recovery in transmitting multiple complex and coherent contextual fragments. Fortunately, state-of-the-art (SOTA) generative AI (GAI) models, have lately emerged as killer applications in many verticals, promising to bring considerable productivity, innovation, and economic value to the real-world services as diverse as image synthesis, text generation, and drug discovery [3]. To be concrete, GAI leverages powerful DL algorithms, such as Transformer and diffusion to automate photorealistic and multimodal AI-generated content (AIGC) in response to user-provided prompts, while its fidelity and accuracy are contingent on adequate pre-training on billions of parameters in large language (e.g., ChatGPT) or image (e.g., Dall-E) models. Most importantly, GAI has immense abilities of context-reasoning and cross-modal content synthesis to generate high-quality and basically correct responses by successfully mimicking human’s thinking and speaking patterns, enabling users feel that they are interacting with a real human-being rather than a dull machine [4, 5, 6, 7]. Such a milestone innovation stimulates us to investigate the potential of applying AIGC into wireless SemCom, which hypothesizes to yield the below benefits. Better SemCom Training Efficiency: Undoubtedly, the growing prosperity of SemCom is inseparable from colossal data resources for semantic model pre-training and background knowledge preparation targeting effective semantic interpretation [2]. To enable better SemCom training efficiency, GAI models are capable of producing vast multimodal content with a certain degree of authenticity and thus should be valuable materials to semantic training and background knowledge. Enhanced Semantic Context-Reasoning: The historical AIGC can be stored online and retrieved easily to provision semantic coding models a better understanding for the context information, thereby offering significant context reasoning and semantic generalization. Furthermore, the creativity of GAI models can offer high-quality and precise content automatically for semantic interpretation in SemCom, thanks to the Prompt Engineer mechanism [8]. This ensures high semantic fidelity even if coding errors occur during joint-source-channel coding (JSCC) or physical signal transmission. Higher Spectrum Utilization: Notice that most of AIGC can be produced via inputting only a few prompts, and the responses can be highly precise and specific if the prompts are well-crafted to align with a task-oriented communication. Hence, by utilizing well-trained GAI models, only several prompts need to be sent, instead of transmitting the whole source information, in each SemCom process to significantly deduct the required bandwidth resources while retaining the original meaning.
@@
Despite many ascendancies offered by the combination of GAI and SemCom, it still encounters several inevitable and thorny challenges that should be tackled before practical implementation. Among them, the paramount issue is how to deal with such considerable computing and storage resources required by these large GAI models. For instance, one of OpenAI’s public large language models, called ChatGPT-3, comprises approximately 175 billion parameters in total [3], and hence it undoubtedly needs to consume colossal computing resources to operate the system. Another problem worth pointing out here lies in the reliability and latency aspects. AIGC is autonomously created that may lead to uncertainty to some extent, while introducing extra delay for data processing as well as data dissemination. To the best of our knowledge, only a few studies have explored the potential of incorporating GAI with SemCom. For example, the authors in [4] and [9] employed different deep GAI networks to enhance the perceptual quality and semantic reliability for SemCom. Similarly, [10] and [11] respectively sought the possibility of using GAI to quantify the semantic importance and adapt the end-to-end transmission rate. However, all of these works consider it from a transceiver design perspective in device-to-device scenarios, while relevant research for the system framework design in general cellular networks is still lacking. To this end, this article focuses on unleashing the full potential of GAI-assisted SemCom network (GAI-SCN) across the cloud-edge-mobile layers, and the main contributions are summarized in a nutshell as follows:
@@
  • We first present the basic concept of SemCom and four major types of GAI technologies, and then provide a comprehensive comparison among traditional communication, SemCom, and GAI-assisted SemCom. Next, we present the potential junctions between SemCom and GAI.
@@
  • We propose a novel GAI-SCN framework that integrates global and local GAI with semantic coding models in a collaborative cloud-edge-mobile design. Afterward, we showcase its viable implementation workflow consisting of three successive stages: Initial Network Preparation Stage, GAI-assisted SemCom Service Provisioning Stage, and Model Synchronization and Update Stage. Moreover, numerical results validate that the proposed framework can save a significant number of transmitted bits while maintaining high-precision semantic delivery compared with two benchmarks.
@@ II-B Typical GAI Technologies
VAEs: The VAE is a likelihood-based generative auto-encoder model, normally comprising of a multi-layer encoder and a symmetric decoder. By taking random training samples with a specific distribution as input, the encoder can regularize its coding distribution to ensure good properties of latent space, while the decoder maps from the latent space to the input space so as to produce new data points [5]. When it comes to the GAI-SCNs, efficient multi-task SemCom is envisioned to be realized with the support of VAEs. DPMs: The DPM is a class of latent variable models, and its generation principle mainly includes two processes of forward diffusion and reverse diffusion. In the forward process, the input content is polluted in steps by introduced Gaussian noise. Afterward, the neural network in the reverse process should be trained to denoise the content blurred with noise, aiming to recover the original data. In comparison of GANs and VAEs, DPMs render remarkable superiority in semantic recovery tasks (e.g., image denoising, inpainting, and super-resolution), producing high-quality content and have better resistance to the risk of noise and interference [6].
@@
Lesson Learned: As introduced, these SOTA GAI technologies are foreseeable to endow SemCom with a high level of semantic interpretation, reproducing and efficient content adjustment by quickly understanding input prompts in a human-thinking manner. However, the localization requirement for semantic models and the cloudification demand for large GAI models are inevitable, and thus it is meaningful to design a realistic GAI-SCN framework in order to yield more benefits from the communication perspective, such as lower consumption of resources, smarter semantic inference, and personalized SemCom services.
@@ II-C Interplay between SemCom and GAI
Keyword Extraction and Goal Identification: Recap that the core of SemCom is meaning delivery, for this purpose, GAI models are capable of extracting some keywords from the input long content, and the corresponding communication goal can be identified in a small amount of text (or the bounding box of objects in image) through excellent context-understanding ability. For instance, Bob uses a mobile device to share his personalized Paris-itinerary plan, detailing his week-long vacation schedule. In this case, several candidate keywords may appear like “One-week long vacation”, “Bob’s travel preferences”, and “Landmarks and recommended food in Paris”, etc., while the goal can be recognized as “Travel plan sharing from Bob”. Accordingly, the global GAI model (e.g., a stable diffusion model) is trained specifically to restore the original content from received keywords and goals given background knowledge about Bob. Moreover, fewer communication resources (including wireless bandwidth and energy) are demanded, while the pressure from stringent latency requirements is relieved.
@@
Lesson Learned: Evidently, the GAI is promising to assist SemCom especially for task-oriented (i.e., meaning delivery-driven), high-capacity, and latency-insensitive services. Notably, although producing AIGC can consume some extra local processing time, the transmission delay is greatly reduced by SemCom in parallel, promising to significantly relieve the latency-cost pressure. Furthermore, we specially compare the system characteristics of GAI-assisted SemCom with traditional communication and SemCom alongside their respective benefits and limitations, as sketched in Fig. 2. Among them, traditional communication is built on bit-based source-and-channel coding following pre-defined and stringent codebooks, while SemCom integrates AI-based semantic encoder and decoder with equivalent background knowledge for accurate semantics-aware JSCC. On this basis, GAI-assisted SemCom takes full advantage of GAI technologies to significantly enhance semantic delivery efficiency and ease resource pressure.
@@
III GAI-Assisted SemCom Network Framework
In this section, we first propose a collaborative GAI-SCN framework, after which its implementation workflow is illustrated step by step.
@@ III-A Hierarchical Structure of GAI-SCN
Consider a SemCom-enabled cellular network scenario as demonstrated in Fig. 3, where there are multiple terminal devices (TDs) of senders and receivers within the coverage of base stations (BSs). Among them, multimodal SemCom services (e.g., text, image, and video) with specific communication goals consecutively arrive at each TD, and each BS acts as the service controller to efficiently schedule and coordinate the goal-oriented SemCom service provisioning. Additionally, a large GAI model (e.g., GPT-4 or Dall-E) is deployed in the cloud to complete computationally intensive tasks, while a small one (e.g., GPT-Neo) is embedded in the TD for coping with local lightweight service demands, such as customized content extraction and text generation services. Specially, we detail the proposed GAI-SCN from the perspective of a three-layer framework design.
@@
Service Provisioning in the Mobile Layer: In the proposed GAI-SCN framework, each TD is equipped with a light-weight GAI model to realize context-aware keyword extraction and goal identification, taking full advantage of its reasoning ability in content understanding. For example, an open-source variant model of ChatGPT, called GPT-Neo, is a family of Transformer-based language models developed by EleutherAI, which has a far smaller size than GPT-3 [12]. Through pre-training on large datasets (like Wikipedia and Common Crawl) and fine-tuning on user data (to be more personalized and customized), GPT-Neo is capable to be directly installed into TDs to precisely extract keywords of source information for users and identify the communication goal in only several words, by which fewer bits and smaller latency are consumed on the transmitter side for data transmission. As for the TDs on the receiver side, semantic decoders are deployed to recover the delivered meaning from obtained bits, after which the GPT-Neo can be further used for semantic calibration or language translation, etc. JSCC Process in the Edge Layer: By enabling sufficient computing and storage ability in distributed edge servers, semantic encoders are considered to be deployed in the edge layer of GAI-SCN. On one hand, we exploit the JSCC method as it is able to greatly improve resilience and robustness of SemCom against various channel states, especially for the case with severe signal impairment, where the essential processes like pre-training, fine-tuning, and reasoning can be completely handled by edge computing servers. On the other hand, considering mass AIGC generated in the cloud or core network, it generally demands considerable computing resources for content pre-processing and semantic encoding before delivering them to users, which can be also offloaded to edge servers. AIGC Acquisition in the Cloud Layer: Based on the above mobile and edge layer designs, the generation of AIGC becomes an indispensable process for smooth SemCom. In the proposed GAI-SCN, a centralized infrastructure, i.e., the remote cloud server, can support and run large GAI models like Google Bard or Microsoft Bing Chat. All preparation processes related to the AI model itself, such as pre-training and fine-tuning, are accomplished in the cloud. Likewise, the collection, analysis, and reasoning for multi-users’ personal information and preferences can be realized in real time by leveraging the massive computing and storage resources of the cloud servers. Moreover, according to the specific keywords and goals uploaded from sender TDs, these large GAI models can quickly and correctly create the response content of desired modes. The rationale behind this is due to the historical conversation record, i.e., referred to as context, between users and GAI servers, which provides reference information for models’ understanding and inference. Therefore, the original meaning implied in the delivered keywords can be recovered to intact text or image content. Lesson Learned: By adopting the above collaborative GAI-SCN framework, the communication resource utilization and performance (e.g., semantic reasoning and generalization capabilities) can be maximized. Particularly, the cloud layer enables global GAI models to periodically coordinate all up-to-date interaction records uploaded from each user, whose personal communicating preferences can be predicted to provide semantic encoders in the edge layer with accurate and integral AIGC services. In parallel, the edge computing servers ensure high reliability and low latency for semantic processing and dissemination conditions via JSCC to provision a very high-quality user experience. Furthermore, light-weight GAI models deployed in the mobile layer can share users’ personalized knowledge and keep it updated in real time. This guarantees consistency of AI’s understanding for user preferences so as to prevent possible information mismatch between keyword extraction and recovered AIGC.

### v6

@@ Abstract
Semantic communication (SemCom) is expected to be a core paradigm in future communication networks, yielding significant benefits in terms of spectrum resource saving and information interaction efficiency. However, the existing SemCom structure is limited by the lack of context-reasoning ability and background knowledge provisioning, which, therefore, motivates us to seek the potential of incorporating generative artificial intelligence (GAI) technologies with SemCom. Recognizing GAI’s powerful capability in automating and creating valuable, diverse, and personalized multimodal content, this article first highlights the principal characteristics of the combination of GAI and SemCom along with their pertinent benefits and challenges. To tackle these challenges, we further propose a novel GAI-integrated SemCom network (GAI-SCN) framework in a cloud-edge-mobile design. Specifically, by employing global and local GAI models, our GAI-SCN enables multimodal semantic content provisioning, semantic-level joint-source-channel coding, and AIGC acquisition to maximize the efficiency and reliability of semantic reasoning and resource utilization. Afterward, we present a detailed implementation workflow of GAI-SCN, followed by corresponding initial simulations for performance evaluation in comparison with two benchmarks. Finally, we discuss several open issues and offer feasible solutions to unlock the full potential of GAI-SCN.
@@ I Introduction
Recently, semantic communication (SemCom) is popularized as an emerging paradigm that promises to significantly alleviate the scarcity of communication resources in future wireless networks [1]. This is mainly benefited from prosper advancement in deep learning (DL) technologies that can drive semantic encoding and decoding models to achieve efficient and high-quality semantic refinement on desired meaning with low spectrum consumption. Moreover, through equipping both ends of the transceiver with equivalent background knowledge, the implicit meaning in conveyed content can be recovered with ultra-low semantic errors even under harsh channel conditions. However, realizing such superiorities obviously poses a huge demand on data acquisition for constructing background knowledge and pre-training DL-driven semantic models. Meanwhile, considering that the achievable semantic performance is essentially confined by the quality of pre-training data used, existing SemCom systems still lack sufficient context reasoning capabilities, i.e., accurate semantic calibration and recovery in transmitting multiple complex and coherent contextual fragments.
Fortunately, state-of-the-art (SOTA) generative AI (GAI) models, have lately emerged as killer applications in many verticals, promising to bring considerable productivity, innovation, and economic value to the real-world services as diverse as image synthesis, text generation, and drug discovery [2]. To be concrete, GAI leverages powerful DL algorithms, such as Transformer and diffusion to automate photorealistic and multimodal AI-generated content (AIGC) in response to user-provided prompts, while its fidelity and accuracy are contingent on adequate pre-training on billions of parameters in large language (e.g., ChatGPT) or image (e.g., Dall-E) models. Most importantly, GAI has immense abilities of context-reasoning and cross-modal content synthesis to generate high-quality and basically correct responses by successfully mimicking human’s thinking and speaking patterns, enabling users feel that they are interacting with a real human-being rather than a dull machine [3, 4, 5, 6]. Such a milestone innovation stimulates us to investigate the potential of applying AIGC into wireless SemCom, which hypothesizes to yield the below benefits.
Better SemCom Training Efficiency: Undoubtedly, the growing prosperity of SemCom is inseparable from colossal data resources for semantic model pre-training and background knowledge preparation targeting effective semantic interpretation [7]. To enable better SemCom training efficiency, GAI models are capable of producing vast multimodal content with a certain degree of authenticity and thus should be valuable materials to semantic training and background knowledge.
Enhanced Semantic Context-Reasoning: The historical AIGC can be stored online and retrieved easily to provision semantic coding models a better understanding for the context information, thereby offering significant context reasoning and semantic generalization. Furthermore, the creativity of GAI models can offer high-quality and precise content automatically for semantic interpretation in SemCom, thanks to the Prompt Engineer mechanism [8]. This ensures high semantic fidelity even if coding errors occur during joint-source-channel coding (JSCC) or physical signal transmission.
Higher Spectrum Utilization: Notice that most of AIGC can be produced via inputting only a few prompts, and the responses can be highly precise and specific if the prompts are well-crafted to align with a task-oriented communication. Hence, by utilizing well-trained GAI models, only several prompts need to be sent, instead of transmitting the whole source information, in each SemCom process to significantly deduct the required bandwidth resources while retaining the original meaning.
@@
Despite many ascendancies offered by the combination of GAI and SemCom, it still encounters several inevitable and thorny challenges in practical implementation. Among them, the paramount issue is how to deal with such considerable computing and storage resources required by these large GAI models. For instance, OpenAI’s product ChatGPT-3 comprises approximately 175 billion parameters in total [2], and hence it undoubtedly needs to consume colossal computing resources to operate the system. Another problem worth pointing out here lies in the reliability and latency aspects. AIGC is autonomously created that may lead to uncertainty to some extent, while introducing extra delay for data processing as well as data dissemination. To the best of our knowledge, only a few studies have explored the potential of incorporating GAI with SemCom. For example, the authors in [3] and [9] employed different deep GAI networks to enhance the perceptual quality and semantic reliability for SemCom. Similarly, [10] and [11] respectively sought the possibility of using GAI to quantify the semantic importance and adapt the end-to-end transmission rate. However, all of these works consider it from a transceiver design perspective in device-to-device scenarios, while relevant research for the system framework design in general cellular networks is still lacking. To this end, this article focuses on unleashing the full potential of GAI-integrated SemCom network (GAI-SCN) across the cloud-edge-mobile layers, and the main contributions are summarized in a nutshell as follows:
@@
  • We first present the basic concept of SemCom and four major types of GAI technologies, and then provide a comprehensive comparison among traditional communication, SemCom, and GAI-integrated SemCom. Next, we present the potential junctions between SemCom and GAI.
@@
  • We propose a novel GAI-SCN framework that integrates global and local GAI with semantic coding models in a collaborative cloud-edge-mobile design. Afterward, we showcase its viable implementation workflow consisting of three successive stages: Initial Network Preparation Stage, GAI-integrated SemCom Service Provisioning Stage, and Model Synchronization and Update Stage. Moreover, numerical results validate that the proposed framework can save a significant number of transmitted bits while maintaining high-precision semantic delivery compared with two benchmarks.
@@ II-B Typical GAI Technologies
VAEs: The VAE is a likelihood-based generative auto-encoder model, normally comprising of a multi-layer encoder and a symmetric decoder. By taking random training samples with a specific distribution as input, the encoder can regularize its coding distribution to ensure good properties of latent space, while the decoder maps from the latent space to the input space so as to produce new data points [4]. When it comes to the GAI-SCNs, efficient multi-task SemCom is envisioned to be realized with the support of VAEs.
DPMs: The DPM is a class of latent variable models, and its generation principle mainly includes two processes of forward diffusion and reverse diffusion. In the forward process, the input content is polluted in steps by introduced Gaussian noise. In comparison of GANs and VAEs, DPMs render remarkable superiority in semantic recovery tasks (e.g., image denoising, inpainting, and super-resolution), producing high-quality content and have better resistance to the risk of noise and interference [5].
@@ II-C Interplay between SemCom and GAI
Keyword Extraction and Goal Identification: Recap that the core of SemCom is meaning delivery, for this purpose, GAI models are capable of extracting some keywords from the input long content, and the corresponding communication goal can be identified in a small amount of text (or the bounding box of objects in image) through excellent context-understanding ability. Note that the extracted semantic granularity with respect to keywords and goals can be flexibly adjusted according to the importance of semantic information combined with each user’s personal preference, especially in the task-oriented semantic transmission scenario. Accordingly, the global GAI model (e.g., a stable diffusion model) is trained specifically to restore the original content from received keywords and goals given background knowledge. Moreover, fewer communication resources (including wireless bandwidth and energy) are demanded, while the pressure from stringent latency requirements is relieved.
@@
Lesson Learned: The GAI is promising to be integrated with SemCom especially for task-oriented (i.e., meaning delivery-driven) and high-capacity modern entertainment communication services, such as mobile virtual reality delivery and Metaverse. Notably, although producing AIGC can consume some extra local processing time, the transmission delay is greatly reduced by SemCom in parallel, promising to significantly relieve the latency-cost pressure. Furthermore, we specially compare the system characteristics of GAI-integrated SemCom with traditional communication and SemCom alongside their respective benefits and limitations, as sketched in Fig. 2. Among them, traditional communication is built on bit-based source-and-channel coding following pre-defined and stringent codebooks, while SemCom integrates AI-based semantic encoder and decoder with equivalent background knowledge for accurate semantics-aware JSCC. On this basis, GAI-integrated SemCom takes full advantage of GAI technologies to significantly enhance semantic delivery efficiency and ease resource pressure.
@@
III GAI-Integrated SemCom Network Framework
@@ III-A Hierarchical Structure of GAI-SCN
Consider a SemCom-enabled cellular network scenario as demonstrated in Fig. 3, where there are multiple terminal devices (TDs) of senders and receivers within the coverage of base stations (BSs). Among them, multimodal SemCom services (e.g., text, image, and video) with specific communication goals consecutively arrive at each TD, and each BS acts as the service controller to efficiently schedule and coordinate the goal-oriented SemCom service provisioning. Additionally, a large GAI model (e.g., GPT-4 or Dall-E) is deployed in the cloud to complete computationally intensive tasks, while a small one (e.g., GPT-Neo [12]) is embedded in the TD for coping with local lightweight service demands, such as customized content extraction and text generation services.
@@
Service Provisioning in the Mobile Layer: Each TD is first equipped with a light-weight GAI model to realize context-aware keyword extraction and goal identification, and compared to data compression methods, this takes full advantage of the context reasoning ability of GAI to guarantee more efficient and intelligent semantic transmission. Through pre-training on large datasets (like Wikipedia and Common Crawl) and fine-tuning on user data (to be more personalized and customized), GPT-Neo is capable to be directly installed into TDs to precisely extract keywords of source information for users and identify the communication goal in only several words, by which fewer bits and smaller latency are consumed on the transmitter side for data transmission. As for the TDs on the receiver side, semantic decoders are deployed to recover the delivered meaning from obtained bits, after which the GPT-Neo can be further used for semantic calibration or language translation, etc. Note that the GPT-Neo can be flexibly replaced by other SOTA DL-based semantic extraction models, as long as they can efficiently accomplish the task of extracting keywords and goals accurately.
JSCC Process in the Edge Layer: By enabling sufficient computing and storage ability in distributed edge servers, semantic encoders are considered to be deployed in the edge layer of GAI-SCN. On one hand, we exploit the JSCC method as it is able to greatly improve resilience and robustness of SemCom especially against the extreme channel conditions, where the essential processes like pre-training, fine-tuning, and reasoning can be completely handled by edge computing servers. On the other hand, considering the large-capacity AIGC transmission scenario that requires excessive wireless network resources, SemCom can ensure not only the resource savings but also the accurate semantic conveying and recovery for the application-layer counterpart, where related semantic coding tasks can be also offloaded to edge servers.
AIGC Acquisition in the Cloud Layer: Based on the above mobile and edge layer designs, the generation of AIGC becomes an indispensable process for smooth SemCom. In the proposed GAI-SCN, a centralized infrastructure, i.e., the remote cloud server, can support and run large GAI models like Google Bard or Microsoft Bing Chat. All preparation processes related to the AI model itself, such as pre-training and fine-tuning, are accomplished in the cloud. Likewise, the collection, analysis, and reasoning for multi-users’ personal information and preferences can be realized in real time by leveraging the massive computing and storage resources of the cloud servers. Moreover, according to the specific keywords and goals uploaded from sender TDs, these large GAI models can quickly and correctly create the response content of desired modes. The rationale behind this is due to the historical conversation record, i.e., referred to as context, between users and GAI servers, which provides reference information for models’ understanding and inference. Therefore, the original meaning implied in the delivered keywords can be recovered to intact text or image content.
Lesson Learned: By adopting the above collaborative cross-layer GAI-SCN framework, both link-level wireless resource utilization and application-layer performance (e.g., satisfaction of user semantic provisioning) can be maximized. Particularly, JSCC-based SemCom is a very appropriate solution to ensure not only resource savings but also accurate key semantics conveying and recovery for the application-layer counterpart. Note that our proposed GAI-SCN does not assume the common phenomenon of signal transmission inconsistency in SemCom, but instead, its whole design is based on the consideration of how to accurately transmit and recover the core semantics between the source and destination. Besides, although the usage cost of GAI-SCN is higher than traditional communication or general SemCom, the savings in communication resources (e.g., spectrums and communication delay) and the promotion in communication quality (e.g., semantic transmission efficiency and resilience to signal distortion) should be considerable as well. Most importantly, the global GAI is placed in the cloud that is generally computing-resource-unlimited, while semantic coding models are preset as maturely trained in advance that means the related training costs are not counted into the budget.